The most recent financial statements for Williamson, Inc., are shown here (assuming no income taxes):

Income Statement Balance Sheet
  Sales $ 8,600   Assets $ 16,100   Debt $ 6,400
  Costs 5,630   Equity 9,700
    Net income $ 2,970     Total $ 16,100     Total $ 16,100

Assets and costs are proportional to sales. Debt and equity are not. No dividends are paid. Next year’s sales are projected to be $10,578.

What is the external financing needed? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

  External financing needed $   
Answer #1

Growth rate in sales=(10,578-8600)/8600=23%

Sales 10,578
Costs(5630*1.23) 6924.9
Net income 3653.1

Total assets would be=$16100*1.23=$19803

Total equity would be=9700+3653.1

=13353.1

Total assets=Total equity+Total debt

Hence external financingn needed=$19803-13353.1-6400

=$50(Approx).